# Optimal quantum simulation of linear non-unitary dynamics

### v1

@@ 1.1 Discussion and conclusion
1.1.1 Related work
Nearing the completion of this manuscript, we became aware of recent work [JLMY25] that also claims optimal complexity using Schrödingerization, though without a discussion of lower bounds. There, a truncation error like the first integral of Equation 8 is analyzed. However, there does not appear to be an error bound equivalent to the second integral of Equation 8 that justifies using functions $f$ that may not be exactly $e^{-x}$ at any $x\geq 0$.
@@ 1.1.2 Inhomogeneous case
The inhomogeneous contribution $\int_{0}^{t}U_{s}(t)\vec{b}(s)\mathrm{d}s$ is obtained by a straightforward modification of the homogeneous case Theorem 4 in more-or-less the same way as described in prior art [ALL23, ACL23], which we describe for completeness. The integral is discretized, such as by a Riemann sum over $M$ points with stepsize $\delta t/M$ like
@@ 1.1.3 Time-dependent simulation
Our main LCHS results Theorems 1, 2, 3 and 4 are stated in full generality for time-dependent $A(t)$. However, the results used to justify optimal query complexity scaling required time-independent $A$. Generally speaking, the query complexity of LCHS of a time-dependent $A(t)$ will be worse by a logarithmic factor in all parameters. This is inherited from the suboptimal query complexity of general time-dependent Hamiltonian simulation by Hermitian $H(t)$, costing $\mathcal{O}(\alpha_{H}t\frac{\log(\alpha_{H}t/\epsilon)}{\log\log(\alpha_{H}t/\epsilon)})$ [LW18] queries to the block-encoding $\textsc{Be}[H/\alpha_{H}]$, compared to the time-independent complexity of $\mathcal{O}(\alpha_{H}t+\log\frac{1}{\epsilon})$. In some special cases though [MF23], the query complexity of time-dependent Hamiltonian simulation can be reduced to the time-independent case.
@@ 1.1.4 $L$ with negative eigenvalues
As noted in [ALL23], any $L$ that is not positive semi-definite can be shifted by an identity term to satisfy the $L\succeq 0$ requirement of LCHS. In the general time-dependent setting, suppose that $L\succeq-l(t)$, for some a priori known time-dependent function $l(t)$. Then let $L_{\text{shifted}}=L+l(t)\mathcal{I}\succeq 0$, and the rescaled homogeneous solution $\vec{w}(t)=e^{-\int_{0}^{t}l(t)\mathrm{d}s}\vec{u}(t)$ satisfies $\frac{\mathrm{d}}{\mathrm{d}t}\vec{w}(t)=-(L_{\text{shifted}}+H(t))\vec{w}(t)$. Although this affects the success probability of obtaining $\lvert u(t)\rangle$, it does not affect the query complexity of Hamiltonian simulation in the LCHS Theorem 2 as $\mathcal{U}_{\text{shifted}}(s;k)=\mathcal{T}e^{-i\int_{0}^{s}k(L(s^{\prime})+l(s^{\prime})\mathcal{I}+H(s^{\prime})\mathrm{d}s^{\prime}}=e^{-ik\int_{0}^{s}l(s^{\prime})\mathrm{d}s^{\prime}}\mathcal{U}(s;k)$, and computation of the $k$- and $s$- dependent phase factor only affects gate complexity. By this change of variables, we may also simulate with the same query complexity $L(t)$ with negative eigenvalues for short times if the negative semi-definite component $L_{-}(t)$ satisfies $\|L_{-}\|_{L^{1}}=\mathcal{O}(1/t)$, which can be possible even if $L(t)$ instantaneously has negative infinite eigenvalues.
@@ 1.1.5 Optimizing constant factors for different applications
In a common scenario, one assumes access to the block-encodings $\textsc{Be}[L/\alpha_{L}]$ and $\textsc{Be}[H/\alpha_{H}]$. In the time-independent case, the overall query complexity of LCHS to these block-encodings followed by amplitude amplification is $\mathcal{O}(\alpha_{\hat{f}}(\alpha_{L}R+\alpha_{H})t+\log\frac{1}{\epsilon}))$. Hence our focus in this work on minimizing the query complexity prefactor $\alpha_{\hat{f}}\cdot R$ is optimal in the limit of large $R=\mathcal{O}(\log\frac{1}{\epsilon})$. However, there will often be cases where $\epsilon$ is large and the decay term $L$ is small compared to $H$ so $\alpha_{L}\ll\alpha_{H}$. In this situation, one should instead numerically optimize $\alpha_{\hat{f}}(\alpha_{L}R+\alpha_{H})$. Alternatively, it was noted in [PJKW25] that by writing $kL+H=\frac{\left(k-i\right)A+(\left(k-i\right)A)^{\dagger}}{2}$, one may block-encode $\textsc{Be}[\frac{kH+L}{\sqrt{R^{2}+1}\alpha_{A}}]$ using one query $\textsc{Be}[A/\alpha_{A}]$. The most expensive term in the LCHS, time-evolution by $e^{-i(RL+H)t}$, is then simulated using $\alpha_{A}\sqrt{R^{2}+1}$ queries, which is slightly better than the naive $R+1$ queries obtained by first block-encoding $\textsc{Be}[L/\alpha_{A}]$ and $\textsc{Be}[H/\alpha_{A}]$ from $\textsc{Be}[A/\alpha_{A}]$. Moreover, if the initial state preparation unitary is considerably more expensive than the matrix block-encodings, one may opt to minimize $\alpha_{\hat{f}}$ alone, and in hybrid near-term quantum-classical schemes with limited depth, one should minimize $R$ alone.
@@ 1.1.6 Fast-forwarding
In this work, we assume access to the block-encoding of the $L\succeq 0$ term. However, previous work for Hermitian $A=L$ has also considered the stronger assumption of access to the block-encoding of $\sqrt{L}$ [CS17, GSLW19, ALWZ25, SGAZ24] . Under this assumption, the query complexity for block-encoding $e^{-Lt}$ is known to scale like $\mathcal{O}\left(\sqrt{\alpha_{L}t+\log\frac{1}{\epsilon}}\sqrt{\log\frac{1}{\epsilon}}\right)$ [LC17a]. Unlike the Hamiltonian simulation problem for $e^{-iHt}$, this square-root dependence on $t$ for exponential decay in the square-root access model does not violate no-fast forwarding. A natural question is whether the LCHS approach can exploit this stronger query model to block-encode the general case $e^{-(L+iH)}$ with the same square-root speedup with respect to $L$, for instance using $\mathcal{O}(\sqrt{\alpha_{L}t\log\frac{1}{\epsilon}}+\alpha_{H}t+\log\frac{1}{\epsilon})$ queries.
@@ 1.1.7 Optimal scaling for general differential equations
Though our LCHS is optimal in the task of block-encoding $e^{-At}$, and in the number of state preparation queries, it is not known to be optimal when applied to solving differential equations. In Table 1, we obtain the normalized solution state $\lvert\vec{u}(t)\rangle$ to error $\epsilon^{\prime}$ using $\mathcal{O}(u_{\text{lchs}}\alpha_{A}t\log\frac{u_{\text{lchs}}}{\epsilon^{\prime}})$ queries to the block-encoding of $A/\alpha_{A}$, and it is not known whether the $u_{\text{lchs}}\log u_{\text{lchs}}$ factor is necessary. We note that this same $\frac{1}{\sqrt{p}}\log\frac{1}{p}$ dependence on final state success probability $p$ also appears in different contexts, such as the current best quantum algorithms for ground state preparation [BTK+25, KLB+25], and existing lower bounds are notably loose [LT20, ALWZ25]. Another natural question is whether a query complexity similar to LCHS can be achieved under more general stability conditions.
@@ 1.1.8 Even better kernel functions
The comparison to the QSVT lower bound indicates that a further constant factor speedup of our LCHS by up to $\approx 3$ may still be possible. Infinitely many other kernel functions beyond those studied in this work satisfy the approximate LCHS conditions Theorem 1, and as shown in the improvement between our examples of $\hat{f}_{2}$ and $\hat{f}_{j,y}$, exploration over the space of kernel functions could narrow this gap. Our LCHS also automatically implies complexity improvements when applied to the more general problem of eigenvalue transformations [ACLY24] for functions beyond $e^{-At}$, and one could hope for further asymptotic improvements by tailoring the kernel functions to each desired function.
@@ 2 Approximate LCHS
In this section, we prove our main mathematical result Theorem 1 that kernel functions that satisfy certain generic analytic conditions lead to a LCHS
@@
Proof of Theorem 1.
By Lemma 2, $\mathcal{U}(t;z)$ is an entire function with respect to $z$. Combined with the definition of $\hat{f}(z)$, the only pole in $\hat{f}(z)\mathcal{U}(t;z)$ on the strip $S_{[-y_{0},0]}$ is at z=iz=-i. Let us choose $\Gamma$ to be a positively-oriented rectangular contour
@@
By definition, $\lim_{x\rightarrow\pm\infty}\hat{f}(x+iy)=0$ for $y\in[-y_{0},0]$. By Lemma 3, $\|\mathcal{U}(t;z)\|\leq 1$ on these segments. Hence, the integrals over $\Gamma_{\text{right}}$ and $\Gamma_{\text{left}}$ vanish as follows.
$\left\|\lim_{x\rightarrow\pm\infty}\int_{0}^{y_{0}}\hat{f}(x+iy)\mathcal{U}(t;x+iy)\mathrm{d}y\right\|$ $\leq\lim_{x\rightarrow\pm\infty}\int_{0}^{y_{0}}|\hat{f}(x+iy)|\|\mathcal{U}(t;x+iy)\|\mathrm{d}y$
$\leq\lim_{x\rightarrow\pm\infty}\int_{0}^{y_{0}}|\hat{f}(x+iy)|\mathrm{d}y$
$\leq\lim_{x\rightarrow\pm\infty}y_{0}\max_{y\in[0,y_{0}]}|\hat{f}(x+iy)|=0.$ (21)
@@
As $\mathcal{U}(t;k)$ is unitary for real $k$, the spectral norm $\|\mathcal{U}(t;k)\|=1$ in the first integral. As $L\succeq 0$Lemma 3 implies that the spectral norm $\|\mathcal{U}(t;k-iy_{0})\|\leq 1$ for $y_{0}\geq 0$ in the second integral. ∎
@@
For all $s\in[0,t]$, let $H(s)$ and $L(s)=L_{+}(s)-L_{-}(s)$, be Hermitian matrices, where $L_{+}$ and $L_{-}$ are the positive semidefinite and negative parts of $L$ respectively. Then for any $t\geq 0$, $z\in\mathbb{C}$,

### v2

@@ 1.1 Discussion and conclusion
1.1.1 Schrödingerization
Schrödingerization [JLY24] solves Equation˜1 under the same conditions as LCHS by a dilation to a higher-dimensional partial differential equation. This perspective is almost equivalent to LCHS, which we elucidate in Appendix˜A. As both in prior art rely on the exact decay equality Equation˜4, both benefit in the same way from the optimal kernel functions of this work. The difference lies in details of how the block-encoding of $e^{-At}$ is implemented, and we demonstrate that the implementation of Schrödingerization in prior art reduces to an asymmetric block-encoding where the magnitude of kernel function coefficients by Prep and $\overline{\textsc{Prep}}$ in Theorem˜4 are not equal. In Appendix˜A, we prove that this asymmetric choice leads to a strictly worse block-encoding normalization compared to the symmetric choice of the LCHS approach. Other kernel functions have also been proposed in the context of Schrödingerization. These are based on smooth extensions to $x<0$ from exact exponential decay for $x\geq 0$ by polynomial interpolation [JLM25] or infinitely-differentiable mollifiers like $\eta(x)=e^{1/(x^{2}-1)}$ [JLMY25] with sub-exponential decay $\hat{\eta}(k)=\Theta(e^{-\sqrt{|k|}})$ [Joh15]. However, these are all subject to no-go results [ACL23, HA25] and cannot decay exponentially fast.
@@ 1.1.2 Inhomogeneous case
The inhomogeneous contribution $\int_{0}^{t}U_{s}(t)\vec{b}(s)\mathrm{d}s$ is obtained by a straightforward modification of the homogeneous case Theorem˜4 in more-or-less the same way as described in prior art [ALL23, ACL23], which we describe for completeness. The integral is discretized, such as by a Riemann sum over $M$ points with stepsize $\delta=t/M$ like
@@ 1.1.3 Time-dependent simulation
Our main LCHS results Theorems˜1, 2, 3 and 4 are stated in full generality for time-dependent $A(t)$. However, the results used to justify optimal query complexity scaling required time-independent $A$. Generally speaking, the query complexity of LCHS of a time-dependent $A(t)$ will be worse by a logarithmic factor in all parameters. This is inherited from the suboptimal query complexity of general time-dependent Hamiltonian simulation by Hermitian $H(t)$, costing $\mathcal{O}(\alpha_{H}t\frac{\log(\alpha_{H}t/\epsilon)}{\log\log(\alpha_{H}t/\epsilon)})$ [LW18] queries to the block-encoding $\textsc{Be}[H/\alpha_{H}]$, compared to the time-independent complexity of $\mathcal{O}(\alpha_{H}t+\log\frac{1}{\epsilon})$. In some special cases though [MF23], the query complexity of time-dependent Hamiltonian simulation can be reduced to the time-independent case.
@@ 1.1.4 $L$ with negative eigenvalues
As noted in [ALL23], any $L$ that is not positive semi-definite can be shifted by an identity term to satisfy the $L\succeq 0$ requirement of LCHS. In the general time-dependent setting, suppose that $L\succeq-l(t)$, for some a priori known time-dependent function $l(t)$. Then let $L_{\text{shifted}}=L+l(t)\mathcal{I}\succeq 0$, and the rescaled homogeneous solution $\vec{w}(t)=e^{-\int_{0}^{t}l(t)\mathrm{d}s}\vec{u}(t)$ satisfies $\frac{\mathrm{d}}{\mathrm{d}t}\vec{w}(t)=-(L_{\text{shifted}}+H(t))\vec{w}(t)$. Although this affects the success probability of obtaining $\lvert u(t)\rangle$, it does not affect the query complexity of Hamiltonian simulation in the LCHS Theorem˜2 as $\mathcal{U}_{\text{shifted}}(s;k)=\mathcal{T}e^{-i\int_{0}^{s}k(L(s^{\prime})+l(s^{\prime})\mathcal{I}+H(s^{\prime})\mathrm{d}s^{\prime}}=e^{-ik\int_{0}^{s}l(s^{\prime})\mathrm{d}s^{\prime}}\mathcal{U}(s;k)$, and computation of the $k$- and $s$-dependent phase factor only affects gate complexity. By this change of variables, we may also simulate with the same query complexity $L(t)$ with negative eigenvalues for short times if the negative semi-definite component $L_{-}(t)$ satisfies $\|L_{-}\|_{L^{1}}=\mathcal{O}(1)$, which can be possible even if $L(t)$ instantaneously has negative infinite eigenvalues – a similar observation is made in concurrent work [HA25] – see Corollary˜2 for a more general discussion.
@@ 1.1.5 Optimizing constant factors for different applications
In a common scenario, one assumes access to the block-encodings $\textsc{Be}[L/\alpha_{L}]$ and $\textsc{Be}[H/\alpha_{H}]$. In the time-independent case, the overall query complexity of LCHS to these block-encodings followed by amplitude amplification is $\mathcal{O}(\alpha_{\hat{f}}((\alpha_{L}R+\alpha_{H})t+\log\frac{1}{\epsilon}))$. Hence our focus in this work on minimizing the query complexity prefactor $\alpha_{\hat{f}}\cdot R$ is optimal in the limit of large $R=\mathcal{O}(\log\frac{1}{\epsilon})$. However, there will often be cases where $\epsilon$ is large and the decay term $L$ is small compared to $H$ so $\alpha_{L}\ll\alpha_{H}$. In this situation, one should instead numerically optimize $\alpha_{\hat{f}}(\alpha_{L}R+\alpha_{H})$. Alternatively, it was noted in [PJKW25] that by writing $kL+H=\frac{\left(k-i\right)A+(\left(k-i\right)A)^{\dagger}}{2}$, one may block-encode $\textsc{Be}[\frac{kL+H}{\sqrt{R^{2}+1}\alpha_{A}}]$ using one query $\textsc{Be}[A/\alpha_{A}]$. The most expensive term in the LCHS, time-evolution by $e^{-i(RL+H)t}$, is then simulated using $\alpha_{A}\sqrt{R^{2}+1}$ queries, which is slightly better than the naive $R+1$ queries obtained by first block-encoding $\textsc{Be}[L/\alpha_{A}]$ and $\textsc{Be}[H/\alpha_{A}]$ from $\textsc{Be}[A/\alpha_{A}]$. Moreover, if the initial state preparation unitary is considerably more expensive than the matrix block-encodings, one may opt to minimize $\alpha_{\hat{f}}$ alone, and in hybrid near-term quantum-classical schemes with limited depth, one should minimize $R$ alone.
@@ 1.1.6 Fast-forwarding
In this work, we assume access to the block-encoding of the $L\succeq 0$ term. However, previous work for Hermitian $A=L$ has also considered the stronger assumption of access to the block-encoding of $\sqrt{L}$ [CS17, GSLW19, ALWZ25, SGAZ24] . Under this assumption, the query complexity for block-encoding $e^{-Lt}$ is known to scale like $\mathcal{O}\left(\sqrt{\alpha_{L}t+\log\frac{1}{\epsilon}}\sqrt{\log\frac{1}{\epsilon}}\right)$ [LC17a]. Unlike the Hamiltonian simulation problem for $e^{-iHt}$, this square-root dependence on $t$ for exponential decay in the square-root access model does not violate no-fast-forwarding. A natural question is whether the LCHS approach can exploit this stronger query model to block-encode the general case $e^{-(L+iH)t}$ with the same square-root speedup with respect to $L$, for instance using $\mathcal{O}(\sqrt{\alpha_{L}t\log\frac{1}{\epsilon}}+\alpha_{H}t+\log\frac{1}{\epsilon})$ queries.
@@ 1.1.7 Optimal scaling for general differential equations
Although our LCHS is optimal in the task of block-encoding $e^{-At}$, and in the number of state preparation queries, it is not known to be optimal when applied to solving differential equations. In Table˜1, we obtain the normalized solution state $\lvert\vec{u}(t)\rangle$ to error $\epsilon^{\prime}$ using $\mathcal{O}(u_{\text{lchs}}\alpha_{A}t\log\frac{u_{\text{lchs}}}{\epsilon^{\prime}})$ queries to the block-encoding of $A/\alpha_{A}$, and it is not known whether the $u_{\text{lchs}}\log u_{\text{lchs}}$ factor is necessary. We note that this same $\frac{1}{\sqrt{p}}\log\frac{1}{p}$ dependence on final state success probability $p$ also appears in different contexts, such as the current best quantum algorithms for ground-state preparation [BTK+25, KLB+25], and existing lower bounds are notably loose [LT20, ALWZ25]. Another natural question is whether a query complexity similar to LCHS can be achieved under more general stability conditions.
@@ 1.1.8 Even better kernel functions
The comparison to the QSVT lower bound indicates that a further constant factor speedup of our LCHS by up to $\approx 3$ may still be possible. Infinitely many other kernel functions beyond those studied in this work satisfy the approximate LCHS conditions Theorem˜1, and as shown in the improvement between our examples of $\hat{f}_{2}$ and $\hat{f}_{j,y}$, exploration over the space of kernel functions could narrow this gap. Our LCHS also automatically implies complexity improvements when applied to the more general problem of eigenvalue transformations [ACLY24] for functions beyond $e^{-At}$, and one could hope for further asymptotic improvements by tailoring the kernel functions to each desired function.
@@ 2 Approximate LCHS
In this section, we prove our main mathematical result Theorem˜1 that kernel functions that satisfy certain generic analytic conditions lead to a LCHS
@@
Proof of Theorem˜1.
By Lemma˜3, $\mathcal{U}(t;z)$ is an entire function with respect to $z$. Combined with the definition of $\hat{f}(z)$, the only pole in $\hat{f}(z)\mathcal{U}(t;z)$ on the strip $S_{[-y_{0},0]}$ is at z=iz=-i. Let us choose $\Gamma$ to be a positively-oriented rectangular contour
@@
Note that the equality $\mathcal{U}(t;-i)=U_{0}(t)$ is rigorous, such that through its power series definition, which is everywhere convergent by Lemma˜3.
By definition, $\lim_{x\rightarrow\pm\infty}\hat{f}(x-iy)=0$ for $y\in[0,y_{0}]$. By Lemma˜4, $\|\mathcal{U}(t;z)\|\leq 1$ on these segments. Hence, the integrals over $\Gamma_{\text{right}}$ and $\Gamma_{\text{left}}$ vanish as follows.
$\left\|\lim_{x\rightarrow\pm\infty}\int_{0}^{y_{0}}\hat{f}(x-iy)\mathcal{U}(t;x-iy)\mathrm{d}y\right\|$ $\leq\lim_{x\rightarrow\pm\infty}\int_{0}^{y_{0}}|\hat{f}(x-iy)|\|\mathcal{U}(t;x-iy)\|\mathrm{d}y$
$\leq\lim_{x\rightarrow\pm\infty}\int_{0}^{y_{0}}|\hat{f}(x-iy)|\mathrm{d}y$
$\leq\lim_{x\rightarrow\pm\infty}y_{0}\max_{y\in[0,y_{0}]}|\hat{f}(x-iy)|=0.$ (21)
@@
As $\mathcal{U}(t;k)$ is unitary for real $k$, the spectral norm $\|\mathcal{U}(t;k)\|=1$ in the first integral. As $L\succeq 0$Lemma˜4 implies that the spectral norm $\|\mathcal{U}(t;k-iy_{0})\|\leq 1$ for $y_{0}\geq 0$ in the second integral. ∎
Note that the error bound Theorem˜1 can be looser than expected and can accommodate unstable dynamics. For instance, the following Corollary˜2 characterizes error in terms of the smallest eigenvalue $l_{0}\preceq L$, which we now allow to be negative. A lack of prior knowledge $l_{0}(t)$ is relevant, as we would otherwise first define a shifted $L_{\text{shifted}}=L-l_{0}\succeq 0$ following 1.1.4 and instead simulate $L_{\text{shifted}}+iH$, which gives the same solution but rescaled to have the best-possible final state norm $\vec{u}_{\text{shifted}}(t)=e^{\int_{0}^{t}l(s)\mathrm{d}s}\vec{u}(t)$.
Corollary 2 (Approximate LCHS with smallest $L$ eigenvalue dependence.).
Under the same conditions as Theorem˜1, except that $L\succeq l_{0}$ for any $l_{0}:[0,t]\rightarrow\mathbb{R}$, the LCHS $O_{R}(t)$ of Equation˜7 satisfies
$\|O_{R}(t)-U_{0}(t)\|$ $\leq\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}\backslash[-R,R]}|\hat{f}(k)|\mathrm{d}k+\frac{e^{-y_{0}\int_{0}^{t}l_{0}(s)\mathrm{d}s}}{\sqrt{2\pi}}\int_{\mathbb{R}}|\hat{f}(k-iy_{0})|\mathrm{d}k.$ (24)
Proof.
The proof identical to that of Theorem˜1 up to Equation˜23. As $\mathcal{U}(t;k)$ is unitary for real $k$, the spectral norm $\|\mathcal{U}(t;k)\|=1$ and the first integral of Equation˜23 for error of truncation to finite $R$ is unchanged. Let $L_{\text{shifted}}(t)=L(t)-l_{0}(t)\succeq 0$. Observe that
$\mathcal{U}(t;z)=\mathcal{T}e^{-i\int_{0}^{t}z(L(s)+H(s)\mathrm{d}s}=e^{-iz\int_{0}^{t}l_{0}(s)\mathrm{d}s}\mathcal{T}e^{-i\int_{0}^{t}z(L_{\text{shifted}}(s)+H(s)\mathrm{d}s},$ (25)
By Lemma˜4, $\|\mathcal{T}e^{-i\int_{0}^{t}(k-iy_{0})(L_{\text{shifted}}(s)+H(s)\mathrm{d}s}\|\leq 1$ for any $y_{0}\geq 0$. Hence we substitute the following bound into the second integral of Equation˜23
$\|\mathcal{U}(t;x-iy_{0})\|\leq|e^{-i(x-iy_{0})\int_{0}^{t}l_{0}(s)\mathrm{d}s}|=e^{-y_{0}\int_{0}^{t}l_{0}(s)\mathrm{d}s}.$ (26)
Corollary˜2 indicates that the error contribution of unstable dynamics where $l_{0}(s)<0$ is canceled by that of stable dynamics where $l_{0}(s)\geq 0$. However, this does not provide an efficient algorithm for simulating arbitrary unstable dynamics, such as by padding $A(s)$ to be identity $\mathcal{I}$ for $s\in(t,T]$ so that $\int_{0}^{T}l(s)\mathrm{d}s=\int_{0}^{t}l(s)\mathrm{d}s+(T-t)=0$, as any possible padding with strictly stable dynamics reduces the final state norm by an exponential factor $e^{-(T-t)}=e^{\int_{0}^{t}l(s)\mathrm{d}s}$.
@@
For all $s\in[0,t]$, let $H(s)$ and $L(s)=L_{+}(s)-L_{-}(s)$, be Hermitian matrices, where $L_{+}$ and $L_{-}$ are the positive semi-definite and negative parts of $L$ respectively. Then for any $t\geq 0$, $z\in\mathbb{C}$,